OpenStudy (anonymous):

Help with differentials Consider the function and express the relationship between a small change in x and the corresponding change in y in the form dy=f'(x)dx f(x)=2x+1

OpenStudy (amistre64):

im curious how that is not just a simple case of taking a derivative

OpenStudy (amistre64):

the rate of change of a line is the slope of the line .....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm having trouble understanding what to do Is the problem. It may not be super difficult I just don't know what to do.

OpenStudy (amistre64):

take the derivative of 2x+1, im assuming youve already been taught how to take a derivative. If not, then I assume you already know that the slope of a line defines the change in y per the change in x

OpenStudy (anonymous):

yes I know both those things. The derivative and slope is just 2

OpenStudy (amistre64):

then dy = 2 dx is how it wants you to express it i believe

OpenStudy (anonymous):

o that's all there is to it? lol. I was thinking to myself, they don't give a dx but you're correct that is all they want, thank you.

OpenStudy (amistre64):

:) for a more complete run thru \[y=2x+1\] \[\frac d{dx}(y=2x+1)\] \[\frac d{dx}(y)=\frac d{dx}(2x+1)\] \[\frac {dy}{dx}=\frac {dx}{dx}2\] \[\frac {dy}{dx}=2\] \[dy=2~dx\]

OpenStudy (anonymous):

ok, thank you